OpenStudy (anonymous):

Can someone show me the steps to solving this? What is the range of the graph of y = –7(x – 1)2 + 2?

OpenStudy (experimentx):

what do you understand by range??

OpenStudy (anonymous):

Nothing :/ aha.

OpenStudy (anonymous):

So would that be y is equal to or less than 2 or y is greater than or equal to 2?

OpenStudy (anonymous):

Could you explain the steps used to find that answer? I'm trying to understand everything.

OpenStudy (experimentx):

less or equal to 2, guess why??

OpenStudy (anonymous):

huh?

OpenStudy (experimentx):

i thought you wanted to find out why??

OpenStudy (anonymous):

oh yeah I do.

OpenStudy (shayaan_mustafa):

still need help?

OpenStudy (anonymous):

Yes.

OpenStudy (shayaan_mustafa):

don't ever give medal until your problem is solved completely or you satisfy completely. OK.

OpenStudy (shayaan_mustafa):

as you required range of the given function. then you have to separate out the x. now do it your self. I am waiting for your answer until.

OpenStudy (anonymous):

aha ohkay & how do you seperate the x?

OpenStudy (anonymous):

Sorry, math is not my best subject.

OpenStudy (shayaan_mustafa):

your original equation is\[\large y=-7(x-1)^{2}+2\]\[\large y+2=-7(x-1)^{2}\]\[\large \frac{y+2}{-7}=(x-1)^{2}\]or write the equation as\[\large (x-1)^{2}= \frac {y+2}{-7}\]taking square root on both sides.\[\large x-1=\pm \sqrt{\frac{y+2}{-7}}\]\[\large x=1\pm \sqrt{\frac{y+2}{-7}}\]It is done now. you have separated x.

OpenStudy (anonymous):

woah that's confusing aha.

OpenStudy (shayaan_mustafa):

which step is making you confuse? tell me i am here to satisfy you miss.

OpenStudy (anonymous):

The second step. Where it turns into y+2 instead of just y.

OpenStudy (shayaan_mustafa):

oh... sorry sorry. it is y-2. on right side, 2 has plus sign. when it moves on left side then its sign becomes - so change in all steps that it is y-2 not y+2

OpenStudy (shayaan_mustafa):

with modification, y-2 \[\large x=1\pm \sqrt{\frac{y-2}{-7}}\]

OpenStudy (anonymous):

Oh ohkay that makes more sense aha

OpenStudy (shayaan_mustafa):

now you got how we can separate x? should we progress now?

OpenStudy (anonymous):

Yes

OpenStudy (shayaan_mustafa):

ok you know the value in the square root must be always positive otherwise we get imaginary number. for example. \[\large y=\sqrt{2}\]this exists. while\[\large y=\sqrt{-2}\]doesn't exist. do you know this?

OpenStudy (anonymous):

Yes I do

OpenStudy (shayaan_mustafa):

ok so the inner term in the square root must be positive i.e. \[\large \frac{2-y}{7}>0\]is it?

OpenStudy (anonymous):

no

OpenStudy (shayaan_mustafa):

so. if y<2 then the quantity becomes positive. if y=0 then the quantity becomes positive. if y>2 then quantity becomes negative and the square root will have - sign in it. so it will ne undefined. so range would be.\[\large (-\infty,2]\] and this range your function will define.

OpenStudy (anonymous):

Oh ohkay thanks! :D

OpenStudy (shayaan_mustafa):

welcome.